OpenStudy (aftong112901):

Which of the following are identities? Check all that apply. a. cot^2x+1=csc^2x b. tan^2x=1-sec^2x c. sin^2x=1+cos^2x d. sin^2x+cos^2x=1

alones (alones):

Which one you think it is?

OpenStudy (aftong112901):

I honestly do not know but I think D is one of them.

OpenStudy (aftong112901):

@AloneS do you know how to do it?

OpenStudy (maurticewilliams):

i can help let me see

OpenStudy (maurticewilliams):

ok i dont know any of these good luck with ur answer

OpenStudy (pythagoras123):

A is correct. B should be \[\tan ^{2}x+1=\sec ^{2}x\] C should be \[\sin ^{2}x + \cos ^{2}x = 1\] D is correct.